Summary of week of 9/3-9/9/99

Ed Blucher (blucher@hep.uchicago.edu)
Fri, 10 Sep 1999 07:45:47 -0500

Summary of E832 Week 13: Sep 3 - Sep 9 Runs: 14418-14488

Hours of beam used/available : 92.8/98.7 = 94%

Vac00 (week) : 371k

Vac00 (99run) : 4.18 million

Comments: In addition to standard datataking, we completed the following
special runs:
1) Reg out to measure transmission of shadow absorber
2) ke3s with ptkick=100 and reg out
3) no CsI threshold run to study threshold correction

We reduced the intensity request to 4E12 on 9/7.

hours of beam:
Shift used/available comment
---------------------------------------------------------------------
Fri Sep 3 OWL: 2.1+4.1/6.6 special ptkick=100 run
DAY: 4.3/5.3 Needed to reset power on ktev3 tapes
EVE: 3.5/3.5 lots of tev probs

Sat Sep 4 OWL: 6.3/6.5
DAY: 7.5/8.1
EVE: 8.25/8.5

Sun Sep 5 OWL: 6.0/6.25
DAY: 7.1/7.3
EVE: 7.3/7.6 ha scaler problems

Mon Sep 6 OWL: 0/0 cryo problems
DAY: 3.2/3.3 left-bend ground fault
EVE: 6.8/7.3 some targetting problems

Tue Sep 7 OWL: 3.5/3.5 TeV Cryo problems
DAY: 7.5/7.75 reduced intensity request to 4e12
EVE: 3.5/3.5 feeder ground fault

Wed Sep 8 OWL: 0/0 feeder ground fault
DAY: 2.7+0/3.3 feeder gnd fault recov.+reg out run
EVE: 0.8+0/0.8 TeV studies + reg out run

Thu Sep 9 OWL: 1.6+4.1/6.1 no CsI thresh run + regular data
DAY: 0/0 Shutdown
EVE: 2.7/3.5 Shutdown recovery
---------------------------------------------------------------------
TOTAL 7.2(special)+85.6(E832)/98.7(available)
---------------------------------------------------------------------